# Equality Saturation for Optimizing High-Level Julia IR

### v1

@@ 3.3. Rules and Types
One of the defining features of the Julia programming language is multiple dispatch. This feature allows users to define a new method of a function for a specific combination of argument types. The fact that the semantics of a function call are not only decided by syntax but also by the type of their arguments necessitates a way to represent this information in rewrite rules as well. For example, a user might want to rewrite matrix multiplications into a function call to some external BLAS library. This means that a call to Base.:*, a function in Julia’s base library, should be rewritten, but only if its arguments are matrices.
Users can specify type constraints in the rewrite rules by associating a type with an unbound variable. For example, a rule for rewriting a matrix multiplication can be written as
We designed a novel e-class analysis for tracking types in the e-graph. For each e-class, this analysis maintains the most specific type that is compatible with all the e-nodes in that class. During e-graph construction, this type is simply the same as the types of statements in the source Julia IR. When two e-classes with a different associated type are merged, the resulting e-class is determined by the typejoin function. For rewrite rules that introduce a new function call, such as the rewrite of a matrix multiplication into blas_call, we run Julia’s type inference using the types stored in the e-classes. Similarly, if the type associated with an e-class changes, we rerun type inference for all e-classes containing e-nodes that depend on the changed e-class.
@@ 3.4. Constants in the E-Graph
Rewrite rules are usually symbolic in nature. They match an expression and replace that expression with a new one, plugging in any matched variable. Modern equality saturation frameworks such as egg (Willsey et al., 2021) and Metatheory.jl (Cheli, 2021b), which was used in this work, also support dynamic rewrites. When dynamic rewrite rules are applied, the right-hand side is first executed (potentially using additional e-class analysis data) and the original code is rewritten into the result of that execution, rather than into the right-hand side itself.
When literals are part of the e-graph, dynamic rewrites can be used to perform simple constant propagation. In normal Julia code, literals of primitive types as well as composite types with a known size can appear directly in the IR. The values of these literals are embedded in the e-graph. The e-classes containing the corresponding e-nodes of the literals have an inferred type associated with them, but compared to e-classes containing nonliteral e-nodes, the actual value of that type is available as well. We extend rewrite rule type annotations with a special Comptime{T} type that allows to match against the type of the actual values in the e-graph instead of the inferred type stored in the e-class. Using instead of to signify dynamic rewrite rules, a rule such as
will match any addition between two integer literals and insert the result of this addition in the respective e-class.
Dynamic rewrite rules also open opportunities for partial evaluation-like optimization (Muchnick, 1998; Futamura, 1999) beyond the constant propagation capabilities of the Julia compiler. The compiler does not propagate dynamic array arguments as their values could be mutated at any time. With dynamic rewrite rules, a programmer can override this default behavior. To achieve this, we extended the rewrite rule format with rules of the form
This rule introduces an equivalence between the function’s second argument in its e-graph and the constant W. A constant array holding W’s contents is added as an equivalent e-node to that argument’s e-class in the e-graph, to which dynamic rewrite rules can then be applied. An interesting case is, for example, when W is a matrix holding the constant weights of a neural network layer. Computations that only depend on W or other compile-time-known variables can then be partially evaluated.
@@ 3.5. Rewrites in the Presence of Side Effects
We depend on the static effect analysis built in the Julia programming language to decide whether a function call has side effects. This analysis determines several program properties for every statement in the IR of a function. The program property effect_free signals whether a statement is free from externally semantically visible side effects. A limitation of Cranelift’s approach using the CFG skeleton is that there is no way to rewrite instructions with side effects. In Cranelift, this is less of a concern as the code that is rewritten is lower level and typically contains a large amount of simple, low-level instructions, many of which are pure. In high-level Julia code, however, the IR we work on generally is much smaller and consists only of a few calls to high-level functions. Because these functions have more complex behavior than simple, low-level instructions, they also often include behavior with side effects. Still, it should be possible for developers to rewrite function calls that have side effects. For example, a user might want to rewrite an expression containing a function that allocates and returns an array.
To enable this, we introduce a technique called CFG skeleton relaxation. We allow users to create rewrite rules that match against functions with side effects. When such a rule matches an expression in the e-graph and the rewrite is applied by adding the equivalence to the e-graph, all function calls that are part of the matched rule are removed from the CFG skeleton.
However, the ability to ignore the effect analysis and overwrite function calls with side effects is not without risks. Users who are unaware of certain side effects might unknowingly remove them and run into hard-to-debug bugs. We mitigate this problem somewhat by requiring an explicit prefix unsafe in front of the rules that are allowed to remove elements from the CFG skeleton.
Although this raises the bar somewhat to specify and use rewrite rules that unintentionally break the intended semantics of a rewritten piece of code, the risk of using rules incorrectly is obviously not mitigated completely. Semantic rule verification is outside the scope of this work. A rewrite rule might be semantically correct in one domain but incorrect in another, such as reduced precision in deep learning, so correctness is up to the rule author. Bad rules can introduce bugs.
@@ 3.6. Extraction
Example function where equality saturation reveals opportunities for reuse. Based on some condition, the function returns either the trace of matrix A or its determinant. The invocations to the functions tr and det that compute the trace respectively the determinant, can be rewritten as a sum and a product over the eigenvalues stored in eig, which is the variable in which the eigenvalues are already stored after a call to the eigen function. The Julia IR is shown next to the source code, and the e-graph is chosen, which consists of 9 e-classes totaling 11 e-nodes. The nine optimal e-nodes, one for each e-class, for extraction are shaded gray.
@@
When Julia IR is generated from an optimized e-graph and its corresponding CFG skeleton, each statement in the CFG skeleton is considered a root node for e-graph extraction. This means that, in contrast to most work on e-graph-based optimizers, we essentially need to run multiple extractions to recover our program. These extractions are not independent of each other. Take, for example, the code and corresponding IR and e-graph in Figure 4. The function f computes and returns two different values eig and x derived from matrix A. eig stores the eigen vectors and values, and x stores the trace or the determinant of the matrix, depending on control flow. In this example, two rules from a collection of linear algebra identities have been applied:
These give rise to the extra e-nodes in the e-classes containing det and tr. Intuitively, we can see that an optimal extraction might be one in which the result of the call to eigen is reused to compute the determinant and/or trace instead of explicitly calling those functions. Indeed, that extraction is valid since the call to eigen dominates both the call to det and to tr so its result is available at those call sites.
@@
We formulate the problem of finding an optimal extraction as an ILP problem. He et al. (He et al., 2017) propose an ILP description for e-graph extraction that ensures an acyclic extracted graph, which is important for further elaboration to produce linearized IR. We adapt and extend their description to take into account e-node reuse from dominating extractions.
An example E-graph with 5 e-classes. E-node n has 2 child e-classes. E-class c contains e-node n, and another, unnamed e-node. Both e-nodes in e-class c have the same two e-node parents. In this graph, nodes are additionally tagged with annotations that illustrate for e-node n the sets p(n) and chi(n).
@@
Let the e-graph be represented by a set of e-nodes , and a set of e-classes . Let be the set of e-nodes that are contained in e-class , and let be the set of children e-classes of a particular e-node . We associate with each e-node a set of e-nodes defined as:
That is, the set of e-nodes that have a child e-class containing . Figure 5 shows an e-graph in which the different parts have been annotated with this notation.
@@
We represent the statements contained in the CFG skeleton as the set . All of these statements cause a different expression to be extracted from the e-graph, rooted at each . For each statement the statements that dominate are represented by the set . Each statement depends on a number of e-class arguments represented by the set .
@@
Finally, we associate a cost with every e-node such that our optimization problem becomes:888Changes to the formulation by He et al. (He et al., 2017) are highlighted in blue.
@@
Where, for each e-node and statement from the CFG skeleton, we have introduced a binary variable . These variables indicate whether the e-node is selected for the extraction rooted at statement . Furthermore, following the work by He et al. (He et al., 2017), each is a class cycle, a collection of classes that contain e-nodes that form a cycle by depending on other classes in the cycle. The e-nodes in an e-class that participate in a class cycle are represented by the set . For each class in each class cycle , we introduce the binary variables that are used to enforce acyclicity in the extracted graph. is assigned 1 if none of the e-nodes in e-class that participate in the class cycle are selected. For a more complete overview and proof of acyclic extraction we refer to the work by He et al. (He et al., 2017). To take into account the dominance information from the CFG skeleton, we adapted their formulation in a few ways.
We adapt constraint (1) to ensure that an e-node can only be selected if for each child e-class an e-node is also selected, or if an e-node is already selected in a dominating extraction .
To prevent e-nodes from being selected in dominating extractions when they are not needed there, we introduce constraint (2). This constraint prevents an e-node from being selected if there is no e-node selected at the same extraction that depends on that e-node.
@@
Constraint (3) ensures that at least one e-node in each e-class that is referenced by a root in the CFG skeleton is truly selected for extraction. Compared to the original ILP formulation by He et al. (He et al., 2017), we added an extra term to account for the fact that an e-node might have already been selected in a dominating extraction.
@@
The remaining constraints are completely analogous to the constraints by He et al. (He et al., 2017) except for the addition that each constraint is added separately for each possible extraction root .
@@
We implement this ILP formulation in JuMP (Lubin et al., 2023), a domain-specific language embedded in Julia for expressing and solving mathematical optimization problems. For a solver, we resort to HiGHS (Huangfu and Hall, 2018). By default, our cost function currently defaults to 1, reducing the optimization problem to choosing the least number of e-nodes. We allow users to manually alter the cost for calls to particular functions to overwrite this behavior.
@@ 4. Use Case Demonstration
To demonstrate the capabilities of our work, this section presents a number of use cases. This demonstration focuses on our main goal of allowing users and domain experts to express and exploit optimization opportunities in a convenient, flexible way. In other words, our use cases aim to demonstrate that our work can give developers the necessary control over optimizations without them having to become compiler experts and without restricting the flexibility and productivity they have come to expect from the Julia ecosystem. In contrast, our work aims to give developers more flexibility, meaning less restrictions, to write code without incurring a reduction in optimization opportunities.
Our work does not at all aim to squeeze more performance out of the Julia programming language and its compiler and run-time implementation. Indeed, any optimization that can be obtained with a rewrite rule can also be obtained by manually rewriting source code, such as by duplicating source code and then special-casing it. This section will hence not evaluate how much the performance of a piece of code can be improved with rewrite rules.
@@ 4.1. Domain Specific Rewrites
Example of domain-specific code used to apply 2D transformations to a point. The function performs a series of transformations: a horizontal translation over distance dx, a rotation over pi radians, a translation over cosinus squared plus sinus squared minus one, which equals zero, another rotation over pi radians, and a translation over minus dx. With a handful of rules, these can be reduced to the identity transformation.
Figure 6 shows a function that applies a sequence of similarity transformations on a 2D point. On close inspection, application of geometric properties, trigonometric identities, and algebraic simplifications can lead to the reduction of all these transformations to a single identity transform. Indeed, the translation on line 8 can be simplified to p by applying
After this translation has been eliminated, the two rotations by radians similarly end up not transforming the point. Lastly, the first and last rotations now cancel each other out as well. After all the corresponding rewrites, the optimized function simply returns its argument p.
With this example, we also illustrate the necessity of CFG skeleton relaxation. The effect inference in the Julia compiler is unable to reason through the rotation matrix construction in the rotate function on line 2 in Figure 6. Because of this, the rewrite rule author is required to allow CFG skeleton relaxation for those rules where rotate is called by prefixing those rules with unsafe.
@@ 4.2. Aiding Multiple Dispatch
Instead of using rewrites purely for simplifying code, they can also be used to generate code that provides additional information to the Julia compiler’ type system to trigger the execution of better performing function implementations through multiple dispatch. An example in the domain of linear algebra is to automatically wrap certain matrix expressions in a new, more specific type that allows more efficient implementations for subsequent computations. Take, for example, the expression between two regular matrices and . For the case where and are the same matrix, that is, , we know that the result is a symmetric matrix. We can encode this fact by applying the following rewrite rule:
The Symmetric function from the LinearAlgebra.jl package takes the upper triangle of its argument and uses that to efficiently represent a symmetric matrix. More importantly, it returns a value of type Symmetric. Other functions in that linear algebra package, e.g., for solving eigenproblems, have specialized implementations for symmetric matrices. After rewriting with the above rule, those specialized implementations will now be called automatically thanks to the multiple dispatch system knowing that they are invoked on type Symmetric instead of on a more generic matrix type.
Another example is the expression , which is symmetric if B is symmetric as well. Again, this fact can be encoded with the following rule:
Both these examples build on the abstractions of the Julia standard library, but the same concepts can also be applied to operations and types in other, external packages.

### v2

@@ 3.2. Rules and Types
With multiple dispatch, the semantics of a function call are decided not only by syntax (i.e., the function name), but also by the types and counts of their arguments. This necessitates a way to represent this information in rewrite rules. For example, a user might want to rewrite matrix multiplications into a function call to some external BLAS library. This means that a call to Base.:*, a base library function, should be rewritten, but only if its arguments are matrices. Users can specify type constraints in the rewrite rules by associating a type with an unbound variable. For example, a rule for rewriting a matrix multiplication can be written as
E-class analyses (Willsey et al., 2021) offer a framework to associate and propagate lattice information in an e-graph. We designed a novel e-class analysis for tracking types in the e-graph. For each e-class, this analysis maintains the most specific type that is compatible with all the e-nodes in that class. During e-graph construction, this type is simply the same as the types of statements in the Julia IR. When two e-classes with a different associated type are merged, the resulting e-class is determined by the typejoin function. For rewrite rules that introduce a new function call, such as the rewrite of a matrix multiplication into blas_call, we run Julia’s type inference using the types stored in the e-classes. Similarly, if the type associated with an e-class changes, we rerun type inference for all e-classes containing e-nodes that depend on the changed e-class. As an optimization, for a small amount of different types that are joined, we construct a Union type that keeps track of all individual possible types. When the union grows too large, we fall back on a single type that is their most specific joint supertype. For example, if a value of type Integer is introduced in an e-class with associated type Float64, the new associated type becomes Union{Integer, Float64}. If, throughout EqSat, this union grows further, the associated type could become Number. Tracking small unions of specific types instead of joining them immediately into their joint supertype enables additional rewrite rules bound to the more specific types.
Rewrite rules that change the type of a value produced by an expression can cause problems. If the value is used by a function that has no method defined for the new type, no valid code can be produced anymore. Our current implementation detects this when the type information of parent nodes is recomputed. When type inference finds no method for arguments with the new type, an error is thrown. It is the responsibility of the rule author to either ensure that rules only introduce new types that are compatible with all the functions that depend on the value, or that new methods are defined to make it so. In future work, we will implement the necessary support to skip the application of rewrite rules that can result in the production of invalid code.
@@ 3.3. Constants in the E-Graph
Rewrite rules are usually symbolic in nature. They match an expression and replace that expression with a new one, plugging in any matched variable. Modern EqSat frameworks such as egg (Willsey et al., 2021) and Metatheory.jl (Cheli, 2021b), which is used in this work, also support dynamic rewrites. When dynamic rewrite rules are applied, the expression in the right-hand side is first evaluated (potentially using additional e-class analysis data) and the original code is rewritten into the result of that execution, rather than into the literal right-hand side expression itself.
When literals are part of the e-graph, dynamic rewrites can be used to perform simple constant folding. In normal Julia code, literals of primitive types and composite types with known size can appear directly in the IR. The values of these literals are embedded in the e-graph. The e-classes containing the corresponding e-nodes of the literals have an inferred type associated with them, but compared to e-classes containing nonliteral e-nodes, the actual value of that type is available as well. To support constant folding of literals, we extend rewrite rule type annotations with a special Comptime{T} type that only allows to match against literals of type T. Using instead of to signify dynamic rewrite rules, a rule such as will match any addition between two integer literals and insert the constant-folded result of this addition in the respective e-class.
Dynamic rewrite rules also open opportunities for partial evaluation-like optimization (Muchnick, 1998; Futamura, 1999) beyond the constant propagation capabilities of the Julia compiler. The compiler does not propagate dynamic array arguments as their values could be mutated at any time. With dynamic rewrite rules, a programmer can override this default behavior. To achieve this, we extended the rewrite rule format with rules of the form This rule introduces an equivalence between the function’s second argument in its e-graph and the constant W. A constant array holding W’s contents is added as an equivalent e-node to that argument’s e-class in the e-graph, to which dynamic rewrite rules can then be applied. An interesting case is, for example, when W is a matrix holding the constant weights of a neural network layer. Computations that only depend on W or other compile-time-known variables can then be partially evaluated.
@@ 3.4. Rewrites in the Presence of Side Effects
As discussed in Section 3.1, the CFG skeleton is used to ensure that all statements with side effects, such as the println in the example of Fig. 3, are materialized when the final optimized program is extracted.
We leverage Julia’s built-in static effect analysis to determine whether a function call has side effects. This analysis determines several properties for each IR statement. The program property effect_free signals whether a statement is “free from externally semantically visible side effects” (jul, 2025). For our purpose, however, this analysis is often too strict and too imprecise, resulting in some functions not being labeled effect_free, even though they do not have side effects that require them to remain in a program. For example, many functions that allocate memory are not marked effect-free, even if the allocated memory never gets accessed again. As a result, even if rewrite rules subsequently introduce equalities that would make calls to these functions superfluous, they will still be included in the optimized code. In short, there is a mismatch between Julia’s effect analysis and the desired freedom to operate of our rewrite system.
In Cranelift, side effects are less of a concern because the code it rewrites is lower level and typically contains a large amount of simple, low-level instructions, many of which are pure. In high-level Julia code, by contrast, the IR we work on generally is much smaller and consists only of a few calls to high-level functions/methods. Because these functions/methods have more complex behavior than simple, low-level instructions, they also often include behavior with potential side effects according to Julia’s built-in analysis.
To overcome this mismatch when rewriting Julia code, we allow users to prefix rewrite rules with the relaxed keyword. Whenever a relaxed rule matches code, all the function calls that are part of the CFG skeleton and that occur in the left-hand side of the rule are “detached” from the CFG skeleton. When the final optimized code is extracted from the e-graph, detached statements are not forced to be materialized. We call this technique CFG skeleton relaxation.
Technically, it is possible to write relaxed rules that get rid of side effects that are depended upon by other parts of the code, thereby breaking its intended semantics. We note, however, that all practical rewrites that we encountered and that needed relaxation as we will discuss in Section 4, did not introduce problems.
Finally, we note that semantic rule verification is outside the scope of this work. As a rewrite rule might be correct in one domain but incorrect in another, such as reduced precision in deep learning, correctness is up to the rule author.
@@ 3.5. Extraction
Example function where EqSat reveals opportunities for reuse. Based on some condition, the function returns either the trace of matrix A or its determinant. The invocations to the functions tr and det that compute the trace respectively the determinant, can be rewritten as a sum and a product over the eigenvalues stored in eig, which is the variable in which the eigenvalues are already stored after a call to the eigen function. The Julia IR is shown next to the source code, and the e-graph is chosen, which consists of 9 e-classes totaling 11 e-nodes. The nine optimal e-nodes, one for each e-class, for extraction are shaded gray.
@@
When Julia IR is generated from an optimized e-graph and its corresponding CFG skeleton, each statement in the CFG skeleton is considered a root node for e-graph extraction. This means that, in contrast to vanilla egraph extraction, we essentially need to run multiple extractions on the single e-graph to recover our program (Coward et al., 2023a). These extractions are not independent of each other. Take, for example, the code and corresponding IR and e-graph in Fig. 4. The method f computes and returns two different values eig and x derived from matrix A. eig stores the eigen vectors and values, and x stores the trace or the determinant of the matrix, depending on control flow. In this example, two rules from a collection of linear algebra identities have been applied: and . These give rise to the extra e-nodes in the e-classes containing det and tr. Intuitively, we can see that an optimal extraction might be one in which the result of the call to eigen is reused to compute the determinant and/or trace instead of explicitly calling those functions. Indeed, that extraction is valid since the call to eigen dominates both the call to det and to tr, so its result is available at these call sites.
@@
We formulate the problem of finding an optimal extraction as an ILP problem. He et al. (He et al., 2023) propose an ILP description for e-graph extraction that ensures an acyclic extracted graph, which is important for further elaboration to produce linearized IR. We adapt and extend their description to take into account e-node reuse from dominating extractions.
@@
Let the e-graph be represented by a set of e-nodes , and a set of e-classes . Let be the set of e-nodes contained in e-class , and let be the set of children e-classes of a particular e-node . We associate with each e-class a set of e-nodes defined as . That is, the set of parent e-nodes that have a child e-class . Fig. 5 shows an e-graph and CFG skeleton in which the different parts have been annotated with this notation.
@@
We represent the statements contained in the CFG skeleton as the set . All of these statements cause a different expression to be extracted from the e-graph, rooted at each . For each statement , the statements that dominate are represented by the set . Each statement depends on a number of e-class arguments represented by the set .
@@
Finally, we associate a cost with every e-node such that our optimization problem becomes111Changes to the formulation by He et al. (He et al., 2023) are highlighted in blue.
@@
Where, for each e-node and statement from the CFG skeleton, we have introduced a binary variable . These variables indicate whether the e-node is selected for the extraction rooted at statement . Furthermore, following the work by He et al. (He et al., 2023), each is a class cycle, a collection of classes that contain e-nodes that form a cycle by depending on other classes in the cycle. The e-nodes in an e-class that participate in a class cycle are represented by the set . For each class in each class cycle , we introduce binary variables to enforce acyclicity in the extracted graph. is assigned 1 if none of the e-nodes in e-class that participate in the class cycle are selected. For a more complete overview and proof of acyclic extraction, we refer to the work by He et al. (He et al., 2023). To take into account the dominance information from the CFG skeleton, we adapted their formulation in a few ways.
@@
We adapt constraint (3) to ensure that an e-node can only be selected if for each child e-class an e-node is also selected, or if an e-node is already selected in a dominating extraction . To prevent e-nodes from being selected in dominating extractions when they are not needed there, we introduce constraint (4). This constraint prevents an e-node from being selected if there is no e-node selected at the same extraction that depends on that e-node. Constraint (5) ensures that at least one e-node in each e-class that is referenced by a root in the CFG skeleton is truly selected for extraction. Compared to the original ILP formulation by He et al. (He et al., 2023), we added an additional term to account for the fact that an e-node might already have been selected in a dominating extraction.
@@
The other constraints are completely analogous to the constraints by He et al. (He et al., 2023) except for the addition that each constraint is added separately for each possible extraction root .
@@
We implemented this ILP formulation in JuMP (Lubin et al., 2023), a domain-specific language embedded in Julia to express and solve mathematical optimization problems. For a solver, we resorted to HiGHS (Huangfu and Hall, 2018). By default, our cost function currently defaults to 1, reducing the optimization problem to choosing the least number of e-nodes. We allow users to manually alter the cost for calls to particular functions to overwrite this behavior.
@@ 4. Use Case Demonstration
Example of domain-specific code used to apply 2D transformations to a point. The method performs a series of transformations: a horizontal translation over distance dx, a rotation over pi radians, a translation over cosine squared plus sine squared minus one, which equals zero, another rotation over pi radians, and a translation over minus dx.
To demonstrate the capabilities of our work, this section presents a number of use cases, focusing on our main goal of allowing users and domain experts to express and exploit optimization opportunities in a convenient, flexible way. In other words, our use cases aim to demonstrate that our work can give developers the necessary control over optimizations without having to become compiler experts and without restricting the flexibility and productivity they have come to expect from the Julia ecosystem. In contrast, our work aims to give developers more flexibility, meaning less restrictions, to write code without incurring a reduction in optimization opportunities.
@@ 4.1. Domain Specific Rewrites
Fig. 6 shows a method that applies a sequence of similarity transformations on a 2D point. On close inspection, under the assumption that exact floating point results must not be maintained, application of geometric properties, trigonometric identities, and algebraic simplifications can lead to the reduction of all these transformations to a single identity transform. Indeed, the translation on line 8 can be simplified to p by applying the following three rules: , , and . After this translation has been eliminated, the two rotations by radians similarly end up not transforming the point. Lastly, the first and last rotations now cancel each other out as well. After all the corresponding rewrites, the optimized method simply returns its argument p.
This example also illustrates the necessity of CFG skeleton relaxation. The Julia compiler’s effect inference fails to reason through the rotation matrix construction in the rotate function on line 2 in Fig. 6. As a result, the allocation of the rotation matrix is conservatively considered to have side effects, which means that, by default, even if an optimization makes the rotation matrix obsolete, the matrix will still be materialized by the optimized code. The rewrite rule author is hence required to allow CFG skeleton relaxation for rules with rotate by prefixing those rules with relaxed.
@@ 4.2. Aiding Multiple Dispatch
Instead of using rewrites purely for simplifying code, they can also be used to generate code that provides additional information to the Julia compiler’ type system to trigger the execution of better performing methods through multiple dispatch. An example in the domain of linear algebra is to automatically wrap certain matrix expressions in a new, more specific type that allows more efficient implementations for subsequent computations. Take, for example, the expression between two regular matrices and . For the case where and are the same matrix, that is, , we know that the result is a symmetric matrix. We can encode this fact by applying the following rewrite rule: .
The Symmetric function from the LinearAlgebra.jl package takes the upper triangle of its argument and uses that to efficiently represent a symmetric matrix. More importantly, it returns a value of type Symmetric. Other functions in that linear algebra package, e.g., for solving eigenproblems, have specialized methods for symmetric matrices. After rewriting with the above rule, those specialized methods will now be called automatically thanks to the multiple dispatch system knowing that they are invoked on the type Symmetric instead of on a more generic matrix type. Another example is the expression , which is symmetric if B is symmetric as well, as can be encoded with the rule .
Both examples build on the abstractions of the Julia standard library, but the same concepts can also be applied to operations and types in other, external packages.